# Antonio Ramírez-de-Arellano

### v2

@@ Abstract
In the present work, we further study the computational power of virus machines (VMs in short). VMs provide a computing paradigm inspired by the transmission and replication networks of viruses. VMs consist of process units (called hosts) structured by a directed graph whose arcs are called channels and an instruction graph that controls the transmissions of virus objects among hosts. The present work complements our understanding of the computing power of VMs by introducing normal forms; these expressions restrict the features in a given computing model. Some of the features that we restrict in our normal forms include (a) the number of hosts, (b) the number of instructions, and (c) the number of virus objects in each host. After we recall some known results on the computing power of VMs we give our series of normal forms, such as the size of the loops in the network, proving new characterisations of family of sets, such as finite sets, semilinear sets, or recursively enumerable sets ($NRE$).
@@ 1 Introduction
In the present work, we consider some normal forms for virus machines, this computing paradigm was introduced in [1] are unconventional and natural computing models inspired by networks of virus replications and transmissions. More information on unconventional and natural computing is found in [2] and [3], respectively. From [1, 4] it is shown that VMs are Turing complete, that is, they are algorithms capable of general-purpose computations. From such works some VMs for computing classes of (in)finite sets of numbers are also shown. Providing normal forms for VMs allows a more refined or deeper view of their computations: What features and their values for VMs can be increased or decreased to increase or decrease computing power?
@@
Briefly, the idea of a normal form for some computing model is to consider restrictions in the model while maintaining its computer power. That is, considering lower bounds for ingredients in a computing model is a natural direction for investigation. For instance, a well-known normal form in language theory is the Chomsky normal form, CNF in short, from [5]. Instead of having an infinite number of forms to write rules in a grammar for context-free sets, CNF shows that two forms are enough. Normal forms in unconventional or bio-inspired models include spiking neural P systems [6] and cellular automata [7], with recent and optimal results in [8, 9], and a recent survey in [10]. In addition to restriction while maintaining the same power, normal forms can provide frontiers. For instance, by giving some lower bound for some value, further decreasing the value can mean that we only compute a proper subset of problems as before.
This paper contributes the following to the study of virus machines and their computing power. Normal forms, some of which are optimal bounds, for VMs are provided in the following sense: (a) providing characterisations (previously were inclusions) for generating families of finite sets; (b) showing new characterisations for finite sets of numbers using restrictions on the number of required hosts, instructions, or viruses; and (c) new characterisations are also given for singleton sets of numbers and some linear progressions or combinations. The previous work was focused on the restriction of only thre ingredients: hosts, instructions, and number of virueses. In this work, new restrictions are considered: limiting or not the host or instruction graphs to be a tree graph, that is, an acylcic directed graph. Moreover, the instruction-channel graph can also be limited in the sense that one channel can be attached to at most one instruction. We show, for instance, that some VMs with a tree instruction or host graph and with some lower bounds on the number of hosts, instructions, and viruses can only compute finite sets. Lastly, we highlight a new (and better) characterisation of semilinear sets of numbers, in short $SLIN$, with virus machines. Our results on normal forms are then used to ask new questions regarding other normal forms and restrictions on VMs.
The present work is a much extended and revised version of the preliminary report in [11]. For instance, results on the family $SLIN$ is especially interesting from the point of view of theory and applications: $SLIN$ is a class enjoying the fact that it is above the family of finite sets $NFIN$ and below the family of Turing computable sets $NRE,$ with $SLIN$ known to be decidable [12]; the decidability of $SLIN$ helps toward the computational complexity of machines for it; applications can benefit from the decidability of $SLIN$, including formal verification, proof assistant [13].
@@ 2.1 Virus Machines
  • $H=\{h_{1},\dots,h_{p}\}$ is the ordered set of hosts, $h_{out}$ can be either in $H$ or not (for this work, we will suppose always $h_{out}=h_{0}\notin H$, $I=\{i_{1},\dots,i_{q}\}$ the ordered set of instructions.
@@
  • $G_{C}=(E_{H}\cup I,E_{I})$ is a unweighted bipartite graph called channel-instruction graph, where the disjoint partition associated is $\{E_{H}\cup I\}$, although a channel can be attached to more than one instruction, each instruction is attached to only one channel at most.
@@
This paper is focused on the computational power of VMs in the generating mode, that is the numbers that can be generated in all the possible computations, and for that we fix the same notation as in the foundational paper [14] which will be called old ingredients. Let $\mathrm{NVM}(p,q,n)$ be the family of sets of natural numbers generated by virus machines with at most $p$ hosts, $q$ instructions, and $n$ viruses in each host at any instant of the computation. For unbounded restrictions, they are replaced by *.
@@ 2.1.1 Explicit Example
To clarify the behavior of these devices, the generating mode, and their visual representation, let us see an explicit example. Let $\Pi_{ex}$ be the VM of degree $(2,4)$ in the generating mode and visually described in Figure 1. The squares represent the host units with $h_{0}$ being the environment, and the double arrows the arcs of the host graph, for simplicity, if the weight $1$ then the number is omitted in the representation. The blue dots are the instructions units and the vertices in the instruction graph. Lastly, the red dashed lines are the edges of the instruction-channel graph.
@@
The computation works as follows, the initial configuration is $\mathcal{C}_{0}=(2,2,i_{1},0)$. The instruction $i_{1}$ is connected to the channel $(h_{1},h_{0}),$ as the host $h_{1}$ is not empty, there is virus transmission and the highest weight path is followed, that is going again to $i_{1}$ with configuration $\mathcal{C}_{1}=(1,2,i_{1},1)$. The process is analogous and configuration $\mathcal{C}_{2}=(0,2,i_{1},2)$ is reached. From here, as host $h_{1}$ is empty, there is no virus transmission and the least weight path is followed, having the configuration $\mathcal{C}_{3}=(0,2,i_{2},2)$. Instruction $i_{2}$ opens the channel $(h_{2},h_{1})$ so there is virus transmission, but now the highest weight path is not unique, having two possible configurations depending on the instruction that is chosen:
@@ 2.1.2 Formal Verification
One technique to mathematically formal verify the integrity of these devices is by designing invariant formulas which highlight interesting properties of the most relevant loops. For the seek of simplicity, some formal verifications will show the invariant formula and how it works; the mathematical proof of these formulas can be easily obtained with the induction technique. Lastly, we propose a novel technique for formal verification, that is, by fixing the sizes of the loops in the instruction graph.
@@ 2.2 Topology
$(v_{1},\dots,v_{n})$ of vertices such that $(v_{i},v+1)\in E$ for $i=1,\dots,n-1$ and $v_{i}\neq v_{j},$ for $i,j=1,\dots,n,$ and $i\neq j$. Under the same conditions, if $(v_{n},v_{1})\in E$, then the path is called a cycle. A graph without cycles is called a tree. The depth of a tree is the longest path of the tree.
@@
Let $\Pi$ be the virus machine fixed in the statement, setting the instruction graph to $D_{I}=(I,E_{I},w_{I})$, as it is not a rooted tree with root $i_{1}$; then $I(i_{1})\neq I$. Let $\Pi^{\prime}$ be the virus machine of degree $(p,q^{\prime})=|I(i_{1})|)$, defined as $\Pi$ but with a new instruction graph $D_{I(i_{1})}=(I(i_{1}),E_{I(i_{1})},w_{I(i_{1})})$.
@@ 2.3 State-of-the-art
The state-of-the-art is presented in Table 1. The virus machines in the generating mode are Turing Universal; that is, they can generate recursively enumerable sets of numbers ($NRE$[1] for unbounded restrictions. This power is severely reduced when the last ingredient is reduced; more precisely, a semilinear set characterisation ($SLIN$) is proved for $\mathrm{NVM}(*,*,2)$ [14]. From now on, not characterisations but inclusions have been proven, for finite sets ($\mathrm{NFIN}$) they are contained in $\mathrm{NVM}(1,*,*)$ and $\mathrm{NVM}(*,*,1)$ [4]. Finally, the set of power of two numbers is contained in $\mathrm{NVM}(2,7,*)$ [4].
@@
$NRE$ [1] = * * *
$SLIN$ [14] = * * 2
$\mathrm{NFIN}$ [4] $\subseteq$ 1 * *
@@
An interesting and natural question is can we further restrict or provide better lower bounds, for known results about VMs? That is, provide “better” characterisations of finite sets or even other families of sets such as the singleton sets, see, for instance, Table 1. As we focus on finite sets later, let us see the VMs used in [4] to generate finite sets. For $\mathrm{NVM}(1,*,*)$ the VM presented in Figure 2, and for $\mathrm{NVM}(*,*,1)$ the Figure 3. The corresponding lemmas were called (viruses) and (hosts), respectively, and we follow the same notation in this work.
@@ 3 Novel results with old ingredients
Having set the state-of-the-art of VM in generating mode, let us see new results by the boundary of those old ingredients: the hosts, the instructions, and the viruses during the computation. This section will start with finite sets, just to complete the previous results, and then it is organised from weak to stronger power (in the sense of bigger sets), in the end a table summarising all the novel results is presented.
@@ 3.1 Finite sets
Let $F=\{m_{1},\dots,m_{k}\}$ a nonempty finite set of natural numbers that $m_{i}>0$. Then $F$ can be generated by a virus machine of $2$ hosts, $12k+$ 1 instructions, and the $2$ virus in each host at most.
@@
for each $10\leq x\leq m_{i}-$ 1. In particular, $\varphi(m_{i}-1)$ is true, let us suppose that $m_{i}$ is odd, then the following computation is verified:
@@
For $m_{i}$ even the computation is analogous, hence the computation halts in $12m_{i}+$ 1 steps and the number generated is $m_{i}$.
@@
Inclusion is direct by the Lemma 3.1. Let us now focus on inequality; for that, we construct a virus machine from [15], extending the work from [16], which generates the set of all natural numbers except the zero, which verifies the restrictions of the proposition.
@@ 3.2 Singleton Sets
    Let $\Gamma=\{v\}$ be a singleton set of natural number $v\in\mathbb{N}$, then it can be generated by the VM $\Pi_{sing_{1}}$ of degree $(1,1)$ depicted in Figure 6, the initial configuration is $\mathcal{C}_{0}=(1,i_{1},0)$ and in the following configuration, a virus is consumed and replicated by the weight of the arc, that is, $v$, and sent to the environment, leading to the halting configuration $\mathcal{C}_{1}=(0,\#,v)$. Thus, after one transition step, the set generated is $\{v\}$.
@@
    superset-of-or-equals Suppose any VM with only one host and one virus: the host can only be attached to the environment, and let us fix that the weight of that channel is $w\in\mathbb{N}$. Thus, the only number generated is $w$ or none, depending on the instruction graph (if the computation halts or not). Thus, we generate a singleton set.
@@
    superset-of-or-equals With only one instruction, there are two possibilities in the instruction graph:
@@ 3.3 Finite linear progressions
The computing power of virus machines highly depends on the instruction graph, to show this, let us see the following result when we bound by $2$ the amount of instructions. For this, we will fix the following notation:
Let $\mathrm{NLinFIN}=\bigcup_{x\in\mathbb{N}}\bigcup_{n\in\mathbb{N}}\bigcup_{N\in\mathbb{N}}(\{x+n\cdot i:0\leq i\leq N\})\cup\{\ \emptyset\}$, the family of finite linear progressions. The following result holds.
@@
superset-of-or-equals For any $x,n,N\in\mathbb{N}$, let us see that there exists a VM $\Pi_{Lin}$ of degree $(2,2)$ that generates the set $\{x+n\cdot i:0\leq i\leq N\}$. The virus machine can be depicted in Figure 7.
@@ 3.4 Finite linear combinations
    Let $\Pi_{1}$ be the virus machine of degree $(3,3)$ visually presented in Figure 8. Let us suppose that the number generated is $m=w_{1}x^{\prime}+w_{2}y^{\prime}+r$; then, the following computation holds. The initial configuration is $\mathcal{C}_{0}=(N_{1},N_{2},1,\\ i_{1},0),$ from here, one virus is transmitted from host $h_{1}$ to the environment replicated by $w_{1}$, the following instruction is non-deterministically chosen between $i_{1}$ and $i_{2}$, here we choose $i_{1}$. This process is repeated $N_{1}-x^{\prime}-1$ times. Thus we reach the following configuration, $\mathcal{C}_{x^{\prime}-1}=(N_{1}-x^{\prime}+1,N_{2},1,i_{1},w_{1}(x^{\prime}-1)).$ Here we choose instruction $i_{2}$, leading to the configuration $C_{x^{\prime}}=(N_{1}-x^{\prime},N_{2},1,i_{2},w_{1}x^{\prime})$. Now the process is analogous with instruction $i_{2}$ and host $h_{2}$. After $y^{\prime}$ transition steps, we choose instruction $i_{3}$, which will open the channel $(h_{3},h_{0})$, reaching the halting configuration: $C_{x^{\prime}+y^{\prime}+2}=(N_{1}-x^{\prime},N_{2}-y^{\prime},0,\#,w_{1}x^{\prime}+w_{2}y^{\prime}+r),$ that is, after $x^{\prime}+y^{\prime}+2$ steps, the machine halts and sends $w_{1}x^{\prime}+w_{2}y^{\prime}+r$ viruses to the environment.
@@
superset-of-or-equals The idea of this proof is based on the possible loops that can exist in a virus machine with 3 instructions. It is important to note that the way to maximise the amount of numbers that can be generated by an explicit virus machine depends on the amount (or size) of nondeterministic loops in the device.

### v3

@@ Abstract
In the present work, we further study the computational power of virus machines (VMs in short). VMs provide a computing paradigm inspired by the transmission and replication networks of viruses. VMs consist of process units (called hosts) structured by a directed graph whose arcs are called channels and an instruction graph that controls the transmissions of virus objects among hosts. The present work complements our understanding of the computing power of VMs by introducing normal forms; these expressions restrict the features in a given computing model. Some of the features that we restrict in our normal forms include (a) the number of hosts, (b) the number of instructions, and (c) the number of virus objects in each host. After we recall some known results on the computing power of VMs we give our series of normal forms, such as the size of the loops in the network, proving new characterisations for families of number sets, such as finite sets, semilinear sets, or recursively enumerable sets ($NRE$).
@@ 1 Introduction
In the present work, we consider some normal forms for virus machines [VM] (VMs), this computing paradigm develops unconventional and natural computing models inspired by networks of virus replications and transmissions. More information on unconventional and natural computing is found in [unconcomp_handb2021] and [naco_handbook2012], respectively. From [VM, GCRVM] it is shown that VMs are Turing complete, that is, they are algorithms capable of general-purpose computations. From such works some VMs for computing classes of (in)finite sets of numbers are also shown. Providing normal forms for VMs allows a more refined or deeper view of their computations: What features and their values for VMs can be increased or decreased to increase or decrease computing power?
@@
Briefly, the idea of a normal form for some computing model is to consider restrictions in the model while maintaining its computer power. That is, considering lower bounds for ingredients in a computing model is a natural direction for investigation. For instance, a well-known normal form in language theory is the Chomsky normal form, CNF in short, from [chomsky_formgram1959]. Instead of having an infinite number of forms to write rules in a grammar for context-free sets, CNF shows that two forms are enough. Normal forms in unconventional or bio-inspired models include spiking neural P systems [ibarra_snpnorm2007] and cellular automata [brownca2010], with recent and optimal results in [ivan_snpvarnorm2022, peper_normbca2012], and a recent survey in [francis_acmc2022jour]. In addition to restriction while maintaining the same power, normal forms can provide frontiers. For instance, by giving some lower bound for some value, further decreasing the value can mean that we only compute a proper subset of problems as before.
This paper contributes the following to the study of virus machines and their computing power. Normal forms, some of which are optimal bounds, for VMs are provided in the following sense: (a) providing characterisations (previously were inclusions) for generating families of finite sets; (b) showing new characterisations for finite sets of numbers using restrictions on the number of required hosts, instructions, or viruses; and (c) new characterisations are also given for singleton sets of numbers and some linear progressions or combinations. The previous work was focused on the restriction of only three ingredients: hosts, instructions, and number of viruses. In this work, new restrictions are considered: limiting or not the host or instruction graphs to be a tree graph, that is, an acyclic directed graph. Moreover, the instruction-channel graph can also be limited in the sense that one channel can be attached to at most one instruction. We show, for instance, that some VMs with a tree instruction or host graph and with some lower bounds on the number of hosts, instructions, and viruses can only compute finite sets. Lastly, we highlight a new (and better) characterisation of semilinear sets of numbers, in short $SLIN$, with virus machines. Our results on normal forms are then used to ask new questions regarding other normal forms and restrictions on VMs.
The present work is a much extended and revised version of the preliminary report in [tree_or_not_bwmc2024]. For example, findings concerning the family of sets $SLIN$ are particularly noteworthy from both a theoretical and an applied perspective: $SLIN$ is a class enjoying the fact that it is above the family of finite sets $NFIN$ and below the family of Turing computable sets $NRE,$ with $SLIN$ known to be decidable [ginsspan1966]; the decidability of $SLIN$ helps toward the computational complexity of machines for it; applications can benefit from the decidability of $SLIN$, including formal verification, proof assistant [nipkow2008linear].
@@ 2.1 Virus Machines
  • $H=\{h_{1},\dots,h_{p}\}$ is the ordered set of hosts, $h_{out}$ can be either in $H$ or not (for this work, we will suppose always $h_{out}=h_{0}\notin H$, $I=\{i_{1},\dots,i_{q}\}$ the ordered set of instructions).
@@
  • $G_{C}=(E_{H}\cup I,E_{I})$ is an unweighted bipartite graph called channel-instruction graph, where the disjoint partition associated is $\{E_{H}\cup I\}$, although a channel can be attached to more than one instruction, each instruction is attached to only one channel at most.
@@
This paper is focused on the computational power of VMs in the generating mode, that is the numbers that can be generated in all the possible computations, and for that we fix the same notation as in the foundational paper [computing-VM] which will be called old ingredients. Let $\mathrm{NVM}(p,q,n)$ be the family of sets of natural numbers generated by virus machines with at most $p$ hosts, $q$ instructions, and $n$ viruses in each host at any instant of the computation. For unbounded restrictions, they are replaced by *.
@@ 2.1.1 Explicit Example
To clarify the behavior of these devices, the generating mode, and their visual representation, let us see an explicit example. Let $\Pi_{ex}$ be the VM of degree $(2,4)$ in the generating mode and visually described in Figure 1. The squares represent the host units with $h_{0}$ being the environment, and the double arrows the arcs of the host graph. For simplicity, if the weight is $1$, then the number is omitted in the representation. The blue dots are the instructions units and the vertices in the instruction graph. Lastly, the red dashed lines are the edges of the instruction-channel graph.
@@
The computation works as follows; the initial configuration is $\mathcal{C}_{0}=(2,2,i_{1},0)$. The instruction $i_{1}$ is connected to the channel $(h_{1},h_{0})$, as the host $h_{1}$ is not empty, then there is virus transmission and the highest weight path is followed, that is, it goes again to $i_{1}$ with configuration $\mathcal{C}_{1}=(1,2,i_{1},1)$. The process is analogous and the configuration $\mathcal{C}_{2}=(0,2,i_{1},2)$ is reached. From here, since the host $h_{1}$ is empty, there is no virus transmission and the least weight path is followed, having the configuration $\mathcal{C}_{3}=(0,2,i_{2},2)$. Instruction $i_{2}$ opens the channel $(h_{2},h_{1})$ so there is virus transmission, but now the highest weight path is not unique, having two possible configurations depending on the instruction that is chosen:
@@ 2.1.2 Formal Verification
One technique to mathematically verify the integrity of these devices is by designing invariant formulas which highlight interesting properties of the most relevant loops. For simplicity, some formal verifications will show the invariant formula and how it works; the mathematical proof of these formulas can be easily obtained with the induction technique. Lastly, we propose a novel technique for formal verification, that is, by fixing the sizes of the loops in the instruction graph.
@@ 2.2 Topology
$(v_{1},\dots,v_{n})$ of vertices such that $(v_{i},v_{i+1})\in E$ for $i=1,\dots,n-1$ and $v_{i}\neq v_{j},$ for $i,j=1,\dots,n,$ and $i\neq j$. Under the same conditions, if $(v_{n},v_{1})\in E$, then the path is called a cycle. A graph without cycles is called a tree. The depth of a tree is the longest path of the tree.
@@
Let $\Pi$ be the virus machine fixed in the statement, setting the instruction graph to $D_{I}=(I,E_{I},w_{I})$, as it is not a rooted tree with root $i_{1}$; then $I(i_{1})\neq I$. Let $\Pi^{\prime}$ be the virus machine of degree $(p,q^{\prime})$ with $q^{\prime}=|I(i_{1})|)$, defined as $\Pi$ but with a new instruction graph $D_{I(i_{1})}=(I(i_{1}),E_{I(i_{1})},w_{I(i_{1})})$.
@@ 2.3 State-of-the-art
The state-of-the-art is presented in Table 1. The virus machines in the generating mode are Turing Universal; that is, they can generate recursively enumerable sets of numbers (NRENRE[VM] for unbounded restrictions. This power is severely reduced when the last ingredient is reduced; more precisely, a semilinear set characterisation ($SLIN$) is proved for $\mathrm{NVM}(*,*,2)$ [computing-VM]. From now on, not characterisations but inclusions have been proven, for finite sets ($\mathrm{NFIN}$) they are contained in $\mathrm{NVM}(1,*,*)$ and $\mathrm{NVM}(*,*,1)$ [GCRVM]. Finally, the set of power of two numbers is contained in $\mathrm{NVM}(2,7,*)$ [GCRVM].
@@
NRENRE [VM] = * * *
SLINSLIN [computing-VM] = * * 2
$\mathrm{NFIN}$ [GCRVM] $\subseteq$ 1 * *
@@
An interesting and natural question is can we further restrict or provide better lower bounds, for known results about VMs? That is, provide “better” characterisations of finite sets or even other families of sets such as the singleton sets, see, for instance, Table 1. As we focus on finite sets later, let us see the VMs used in [GCRVM] to generate finite sets. For $\mathrm{NVM}(1,*,*)$ the VM presented in Figure 2, and for $\mathrm{NVM}(*,*,1)$ the Figure 3. The corresponding lemmas were called (viruses) and (hosts), respectively, and we follow the same notation in this work.
@@ 3 Novel results with old ingredients
Having set the state-of-the-art of VM in generating mode, let us see new results by the boundary of those old ingredients: the hosts, the instructions, and the viruses during the computation (see Section 2.1). This section will start with finite sets, just to complete the previous results, and then it is organised from weak to stronger power (in the sense of bigger sets), in the end a table summarising all the novel results is presented.
@@ 3.1 Finite sets
Let $F=\{m_{1},\dots,m_{k}\}$ a nonempty finite set of natural numbers with $m_{i}>0$. Then $F$ can be generated by a virus machine of $2$ hosts, $2k+1$ instructions, and $2$ virus in each host at most.
@@
for each $0\leq x\leq m_{i}-1$. In particular, $\varphi(m_{i}-1)$ is true, let us suppose that $m_{i}$ is odd, then the following computation is verified:
@@
For $m_{i}$ even the computation is analogous, hence the computation halts in $2m_{i}+1$ steps and the number generated is $m_{i}$.
@@
Inclusion is direct by the Lemma 3.1. Let us now focus on inequality; for that, we construct a virus machine from [vm_matriz_bwmc2024], extending the work from [vm_matriz2024], which generates the set of all natural numbers except the zero, which verifies the restrictions of the proposition.
@@ 3.2 Singleton Sets
    Let $\Gamma=\{v\}$ be a singleton set of natural number $v\in\mathbb{N}$, then it can be generated by the VM $\Pi_{sing_{1}}$ of degree $(1,1)$ depicted in Figure 6. Here the initial configuration is $\mathcal{C}_{0}=(1,i_{1},0)$ and in the following configuration, a virus is consumed and replicated by the weight of the arc, that is, $v$, and sent to the environment, leading to the halting configuration $\mathcal{C}_{1}=(0,\#,v)$. Thus, after one transition step, the set generated is $\{v\}$.
@@
    Suppose any VM with only one host and one virus: the host can only be attached to the environment, and let us fix that the weight of that channel is $w\in\mathbb{N}$. Thus, the only number generated is $w$ or none, depending on the instruction graph (if the computation halts or not). Thus, we generate a singleton set.
@@
    With only one instruction, there are two possibilities in the instruction graph:
@@ 3.3 Finite linear progressions
The computing power of virus machines highly depends on the instruction graph; to highlight this, let us see the following result when we bound by $2$ the amount of instructions. For this, we will fix the following notation:
Let $\mathrm{NLinFIN}=\bigcup_{x\in\mathbb{N}}\bigcup_{n\in\mathbb{N}}\bigcup_{N\in\mathbb{N}}(\{x+n\cdot i:0\leq i\leq N\})\cup\{\ \emptyset\}$, be the family of finite linear progressions. The following result holds.
@@
For any $x,n,N\in\mathbb{N}$, let us see that there exists a VM $\Pi_{Lin}$ of degree $(2,2)$ that generates the set $\{x+n\cdot i:0\leq i\leq N\}$. The virus machine can be depicted in Figure 7.
@@ 3.4 Finite linear combinations
    Let $\Pi_{1}$ be the virus machine of degree $(3,3)$ visually presented in Figure 8. Let us suppose that the number generated is $m=w_{1}x^{\prime}+w_{2}y^{\prime}+r$; then, the following computation holds. The initial configuration is $\mathcal{C}_{0}=(N_{1},N_{2},1,i_{1},\\ 0),$ from here, one virus is transmitted from host $h_{1}$ to the environment replicated by $w_{1}$, the following instruction is non-deterministically chosen between $i_{1}$ and $i_{2}$, here we choose $i_{1}$. This process is repeated $N_{1}-x^{\prime}-1$ times. Thus we reach the following configuration, $\mathcal{C}_{x^{\prime}-1}=(N_{1}-x^{\prime}+1,N_{2},1,i_{1},w_{1}(x^{\prime}-1)).$ Here we choose instruction $i_{2}$, leading to the configuration $C_{x^{\prime}}=(N_{1}-x^{\prime},N_{2},1,i_{2},w_{1}x^{\prime})$. Now the process is analogous with instruction $i_{2}$ and host $h_{2}$. After $y^{\prime}$ transition steps, we choose instruction $i_{3}$, which will open the channel $(h_{3},h_{0})$, reaching the halting configuration: $C_{x^{\prime}+y^{\prime}+2}=(N_{1}-x^{\prime},N_{2}-y^{\prime},0,\#,w_{1}x^{\prime}+w_{2}y^{\prime}+r),$ that is, after $x^{\prime}+y^{\prime}+2$ steps, the machine halts and sends $w_{1}x^{\prime}+w_{2}y^{\prime}+r$ viruses to the environment.
@@
The idea of this proof is based on the possible loops that can exist in a virus machine with 3 instructions. It is important to note that the way to maximise the amount of numbers that can be generated by an explicit virus machine depends on the amount (or size) of nondeterministic loops in the device.